# Variational formulations of transport phenomena on combinatorial meshes

### v7

@@ Abstract
We develop primal and mixed variational formulations of transport phenomena on cell complexes with simple polytope connectivity. This framework addresses materials with internal structures comprising components of different topological dimensions, where cells of each dimension may possess distinct physical properties. The approach, which we call Combinatorial Mesh Calculus (CMC), extends Forman’s combinatorial differential forms, previously used to formulate strong conservation laws. CMC operates directly on meshes without requiring smooth embeddings, using discrete analogues of the exterior derivative, Hodge star, and co-differential operators. Our mixed formulation achieves computational efficiency through diagonal stiffness matrix A, which admits direct inversion and enable efficient solution strategies. CMC differs from Discrete Exterior Calculus, which requires circumcentric duality and well-centred meshes, and from Finite Element Exterior Calculus, which constructs polynomial spaces on smooth domains. Our framework applies to general cell complexes, including curved cells and irregular meshes without geometric quality constraints. The mathematical development proceeds in parallel between the smooth and discrete settings, establishing correspondences between continuous and discrete operators. Initial boundary value problems are formulated for mass diffusion, heat conduction, charge transport, and fluid flow through porous media. Numerical examples on regular and irregular meshes in two and three dimensions demonstrate agreement with analytical solutions. The framework enables modelling of transport in materials where microstructural topology influences macroscopic behaviour, with applications to polycrystalline materials, composites, and porous media.
@@ 1.2 Overview and contributions
Although the variational derivations conceptually mirror those in vector calculus, we derive them entirely within the language of exterior calculus. The resulting formulations – particularly the mixed one – are related to known results, such as the mixed variational formulation for the Dirichlet problem, a special case of the more general weak Hodge Laplacian problem in Finite Element Exterior Calculus (FEEC) [arnold2018finite, , Chapter 8]. However, our formulation extends the FEEC approach in two key ways: (1) it accommodates mixed boundary conditions (Dirichlet and Neumann simultaneously); (2) while FEEC focuses on the mathematical analysis of Hilbert complexes and their discretisations, we emphasise physical interpretation, discussing the natural domains and physical dimensions of the involved quantities.
@@
Finally, in Section 6, we verify our approach with several steady-state problems across different spatial dimensions and with varying domain and mesh geometries. While our long-term goal is to apply the method to complex microstructures (as in [berbatov2022diffusion, , Section 4]), here we test it on manufactured solutions of continuum problems and report discretisation errors. The method can be viewed both as a discretisation of continuum models and as an intrinsically discrete modelling framework. The accompanying code, available at https://github.com/kipiberbatov/cmc, includes additional steady-state and transient examples, though only steady-state results are reported in this article for clarity.
@@ 2.1 Topological operations
The antisymmetric property is generalised to the graded commutative rule: $\wedge_{q,p}=(-1)^{pq}\wedge_{p,q}$.
@@
    The above equations in a coordinate system $\{x^{q}\}_{q=1}^{D}$ takes the form
@@ 2.2 Metric operations
Consider an oriented orthonormal basis $(e^{1},...,e^{D})$ of $\Omega^{1}M$, $\mathop{\rm vol}\nolimits:=e^{1}\wedge...\wedge e^{D}$ be the volume form. Let $I=(I_{1},...,I_{p})$, $1\leq I_{1}<...<I_{p}\leq D$ and $e_{I}:=e^{I_{1}}\wedge...\wedge e^{I_{p}}$. Let $J$ be the ordered complement of $I$ (with respect to (1,,D)(1,...,D)), and $e^{J}:=e^{J_{1}}\wedge...\wedge e_{J_{D-p}}$. Then $\star_{p}e^{I}=se^{J}$ for some $s\in\{-1,1\}$ (since taking the wedge product gives $e^{1}\wedge...\wedge e^{D}=\mathop{\rm vol}\nolimits$ after reordering, taken into account by $s$), and similarly $\star_{D-p}e^{J}=te^{I}$ for some $t\in\{-1,1\}$. Then
@@ 3 Exterior calculus variational formulations
Phenomenon
Amount
$\mathsf{X}]$
Potential
$\mathsf{Y}]$
Volumetric capacity
$\mathsf{X}\mathsf{Y}^{-1}\mathsf{L}^{-D}]$
Conductivity
$\mathsf{X}\mathsf{Y}^{-1}\mathsf{L}^{2-D}\mathsf{T}^{-1}]$
Constitutive
law
Mass transfer
Mass
$\mathsf{M}]$
Mass density
$\mathsf{M}\mathsf{L}^{-D}]$
1
$[1]$
Mass diffusivity
$\mathsf{L}^{2}\mathsf{T}^{-1}]$
Fick’s law
Heat transfer
Heat energy
$\mathsf{M}\mathsf{L}^{2}\mathsf{T}^{-2}]$
Temperature
$\theta]$
Volumetric heat capacity
$\mathsf{M}\mathsf{L}^{2-D}\mathsf{T}^{-2}\theta^{-1}]$
Thermal conductivity
$\mathsf{M}\mathsf{L}^{4-D}\mathsf{T}^{-3}\theta^{-1}]$
Fourier’s law
Charge transport
Charge
$\mathsf{C}]$
Electric potential
$\mathsf{M}\mathsf{L}^{2}\mathsf{T}^{-2}\mathsf{C}^{-1}]$
Volumetric capacitance
$\mathsf{M}^{-1}\mathsf{L}^{-2-D}\mathsf{T}^{2}\mathsf{C}^{2}]$
Electrical conductivity
$\mathsf{M}^{-1}\mathsf{L}^{-D}\mathsf{T}\mathsf{C}^{2}]$
Ohm’s law
Fluid flow through
porous media
Volume
$\mathsf{L}^{D}]$
Pressure
$\mathsf{M}\mathsf{L}^{2-D}\mathsf{T}^{-2}]$
Compressibility
$\mathsf{M}^{-1}\mathsf{L}^{D-2}\mathsf{T}^{2}]$
Hydraulic conductivity
$\mathsf{M}^{-1}\mathsf{L}^{D}\mathsf{T}]$
Darcy’s law
@@
Discussion 3.18 (Initial condition).
Since we are describing a transient phenomenon, we need an initial condition. We assume the initial potential is prescribed to some initial value
@@
$u(t_{0})$ $u_{0}$ $\mathsf{Y}$ Initial condition Equation 3.37
@@ 3.1 Primal weak formulations of transport phenomena
$\begin{split}\int_{M}w\wedge\frac{\partial Q}{\partial t}&=-\int_{M}(w\wedge d_{D-1}q)+\int_{M}(w\wedge f)\\ &=-\int_{\partial M}\mathop{\rm tr}\nolimits_{\partial M,D-1}(w\wedge q)+\int_{M}(d_{0}w\wedge q)+\int_{M}(w\wedge f)\\ &=-\int_{\Gamma_{\mathbf{N}}}(\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{N}},0}w\wedge\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{N}},D-1}q)+\int_{M}(d_{0}w\wedge(-\star_{1}\widetilde{\kappa}d_{0}u+\star_{D}Q\wedge v))+\int_{M}(w\wedge f)\\ &=-\int_{\Gamma_{\mathbf{N}}}(\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{N}},0}w\wedge g_{\mathbf{N}})-\langle d_{0}w,\widetilde{\kappa}d_{0}u\rangle_{M,1}+\int_{M}(d_{0}w\wedge(\widetilde{\pi}u\wedge v))+\int_{M}(w\wedge f).\end{split}$ (3.39)
@@
Equating the right-hand sides of Equation 3.39 and Equation 3.40, imposing the Dirichlet boundary condition, Equation 3.19, and the initial condition, Equation 3.37, leads to the primal weak formulation presented below.
@@
$F\colon\Omega^{0}M\to\mathbb{R},\quad$ $F(w):=\int_{M}(w\wedge f)\qquad$ $[\mathsf{X}\mathsf{T}^{-1}].$ (3.41e)
@@
$\forall w[\mathsf{Y}]\in\mathop{\rm Ker}\nolimits\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{D}},0},\quad$ $B(w,\frac{\partial u}{\partial t})+A_{\mathcal{D}}(w,u)-A_{\mathcal{A}}(w,u)$ $=F(w)-G(w)\qquad$ $[\mathsf{X}\mathsf{T}^{-1}\mathsf{Y}],$ (3.42a)
@@
$q(t,x)=\begin{cases}(-\star_{1}\widetilde{\kappa}d_{0}u+\widetilde{\pi}u\wedge v)(t,x),&x\notin\Gamma_{\mathbf{N}}\\ g_{\mathbf{N}}(t,x),&x\in\Gamma_{\mathbf{N}}\end{cases},\ t\in I.$ (3.43)
@@
$\forall w[\mathsf{Y}]\in\mathop{\rm Ker}\nolimits\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{D}},0},\quad$ $A_{\mathcal{D}}(w,u)-A_{\mathcal{A}}(w,u)$ $=F(w)-G(w)\qquad$ $[\mathsf{X}\mathsf{T}^{-1}\mathsf{Y}],$ (3.44a)
@@
$q(x)=\begin{cases}(-\star_{1}\widetilde{\kappa}d_{0}u+\widetilde{\pi}u\wedge v)(x),&x\notin\Gamma_{\mathbf{N}}\\ g_{\mathbf{N}}(x),&x\in\Gamma_{\mathbf{N}}\end{cases}.$ (3.45)
@@ 3.2 Mixed weak formulations of transport phenomena
$\begin{split}\langle r,\kappa^{-1}q\rangle_{M,D-1}&=\langle r,\kappa^{-1}q_{\mathcal{D}}\rangle_{M,D-1}+\langle r,\kappa^{-1}q_{\mathcal{A}}\rangle_{M,D-1}\\ &=-\int_{\Gamma_{\mathbf{D}}}(\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{D}},D-1}r\wedge g_{\mathbf{D}})+\langle d_{D-1}r,\tilde{u}\rangle+\langle r,\kappa^{-1}(\star_{D}\pi\widetilde{u}\wedge v)\rangle_{M,D-1}.\end{split}$ (3.49)
@@
Combining the weak version of the constitutive law (Equation 3.49), the weak version of the conservation law (Equation 3.50), the Neumann boundary condition (Equation 3.15), and the initial condition (Equation 3.37), leads to the formulation presented below.
@@
$A\colon\Omega^{D-1}M\times\mathcal{C}^{\infty}(I,\Omega^{D-1}M)\to\mathbb{R},\$$A(r,s):=\langle r,\kappa^{-1}s\rangle_{M,D-1}\$$[\mathsf{X}^{-1}\mathsf{T}\mathsf{Y}],$ (3.51a)
$B_{\mathcal{D}}\colon\Omega^{D}M\times\mathcal{C}^{\infty}(I,\Omega^{D-1}M)\to\mathbb{R},\$$B_{\mathcal{D}}(\widetilde{w},r):=\langle d_{D-1}r,\widetilde{w}\rangle_{M,D}\$$[\mathsf{L}^{-D}],$ (3.51b)
$B_{\mathcal{A}}\colon\Omega^{D}M\times\mathcal{C}^{\infty}(I,\Omega^{D-1}M)\to\mathbb{R},\$$B_{\mathcal{A}}(\widetilde{w},r):=\langle r,\kappa^{-1}(\star_{D}\pi\widetilde{w}\wedge v)\rangle_{M,D-1}\$$[\mathsf{L}^{-D}],$ (3.51c)
$C\colon\Omega^{D}M\times\mathcal{C}^{\infty}(I,\Omega^{D}M)\to\mathbb{R},\$$C(\widetilde{w},\widetilde{u}):=\langle\pi\widetilde{u},\widetilde{w}\rangle_{M,D}\$$[\mathsf{X}\mathsf{L}^{-2D}\mathsf{Y}^{-1}],$ (3.51d)
$G\colon\Omega^{D-1}M\to\mathbb{R},\$$G(r):=\int_{\Gamma_{\mathbf{D}}}(\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{D}},D-1}r\wedge g_{\mathbf{D}})\$$[\mathsf{Y}],$ (3.51e)
$F\colon\Omega^{D}M\to\mathbb{R},\$$F(\widetilde{w}):=\langle f,\widetilde{w}\rangle_{M,D}\$$[\mathsf{X}\mathsf{T}^{-1}\mathsf{L}^{-D}],$ (3.51f)
$\mathtt{flow\_rate}\colon\Omega^{0}M\to\mathbb{R},\$$\mathtt{flow\_rate}(w):=\kappa\,d^{\star}_{D}\star_{0}w+\pi w\wedge v\$$[\mathsf{X}\mathsf{T}^{-1}\mathsf{Y}^{-1}].$ (3.51g)

### v8

@@ Abstract
We develop primal and mixed variational formulations of transport phenomena on cell complexes with simple polytope connectivity. This framework addresses materials with internal structures comprising components of different topological dimensions, where cells of each dimension may possess distinct physical properties. The approach, which we call Combinatorial Mesh Calculus (CMC), extends Forman’s combinatorial differential forms, previously used to formulate strong conservation laws. CMC operates directly on meshes without requiring smooth embeddings, using discrete analogues of the exterior derivative, Hodge star, and co-differential operators. Our mixed formulation leads to a block-diagonal mass-like matrix arising from inner products weighted by material coefficients, enabling efficient local elimination strategies within the mixed system. CMC differs from Discrete Exterior Calculus, which requires circumcentric duality and well-centred meshes, and from Finite Element Exterior Calculus, which constructs polynomial spaces on smooth domains. Our framework applies to general cell complexes, including curved cells and irregular meshes; nonetheless irregularity leads to worse numerical performance. The mathematical development proceeds in parallel between the smooth and discrete settings, establishing correspondences between continuous and discrete operators. Initial boundary value problems are formulated for mass diffusion, heat conduction, charge transport, and fluid flow through porous media. Numerical examples on regular and irregular meshes in two and three dimensions demonstrate agreement with analytical solutions. The framework enables modelling of transport in materials where microstructural topology influences macroscopic behaviour, with applications to polycrystalline materials, composites, and porous media.
@@ 1.2 Overview and contributions
Although the variational derivations conceptually mirror those in vector calculus, we derive them entirely within the language of exterior calculus. The resulting formulations – particularly the mixed one – are related to known results, such as the mixed variational formulation for the Dirichlet problem, a special case of the more general weak Hodge Laplacian problem in Finite Element Exterior Calculus (FEEC) [arnold2018finite, , Chapter 8]. While FEEC naturally accommodates essential and natural boundary conditions, our emphasis is on their explicit interpretation in terms of primal and dual combinatorial forms and associated physical quantities.
While the algebraic structure of the present framework is closely related to Discrete Exterior Calculus and Finite Element Exterior Calculus, the intent of this work is not to discretise a continuum problem posed on a smooth domain. Instead, physical balance laws are formulated directly on a given cell complex, which is treated as a primary representation of the system. Cell measures and material parameters are introduced subsequently to encode geometry and physics. The framework should therefore be interpreted as a native discrete physical model rather than as a numerical approximation of a continuous one.
@@
Finally, in Section 6, we verify our approach with several steady-state problems across different spatial dimensions and with varying domain and mesh geometries. While our long-term goal is to apply the method to complex microstructures (as in [berbatov2022diffusion, , Section 4]), here we test it on manufactured solutions of continuum problems and report discretisation errors. The method can be viewed both as a discretisation of continuum models and as an intrinsically discrete modelling framework. The accompanying code berbatov2026cmc , https://github.com/kipiberbatov/cmc, includes additional steady-state and transient examples, though only steady-state results are reported in this article for clarity.
@@ 2.1 Topological operations
The antisymmetric property is generalised to the graded-commutative rule: $\wedge_{q,p}=(-1)^{pq}\wedge_{p,q}$.
@@
The above equation in a coordinate system $\{x^{q}\}_{q=1}^{D}$ takes the form
@@ 2.2 Metric operations
Consider an oriented orthonormal basis $(e^{1},...,e^{D})$ of $\Omega^{1}M$, $\mathop{\rm vol}\nolimits:=e^{1}\wedge...\wedge e^{D}$ be the volume form. Let $I=(I_{1},...,I_{p})$, $1\leq I_{1}<...<I_{p}\leq D$ and $e_{I}:=e^{I_{1}}\wedge...\wedge e^{I_{p}}$. Let $J$ be the ordered complement of $I$ (with respect to (1,,D)(1,...,D)), and $e^{J}:=e^{J_{1}}\wedge...\wedge e^{J_{D-p}}$. Then $\star_{p}e^{I}=se^{J}$ for some $s\in\{-1,1\}$ (since taking the wedge product gives $e^{1}\wedge...\wedge e^{D}=\mathop{\rm vol}\nolimits$ after reordering, taken into account by $s$), and similarly $\star_{D-p}e^{J}=te^{I}$ for some $t\in\{-1,1\}$. Then
@@ 3 Exterior calculus variational formulations
For simplicity, assume that the volumetric capacity is time-independent. Then,
$Q(t)=Q(t_{0})+\int_{t_{0}}^{t}\frac{\partial Q}{\partial\tau}\,d\tau=Q(t_{0})+\int_{t_{0}}^{t}\star_{0}\widetilde{\pi}\left(\frac{\partial u}{\partial\tau}\,d\tau\right)=Q(t_{0})-\star_{0}(\widetilde{\pi}u)(t_{0})+\star_{0}(\widetilde{\pi}u)(t)=\star_{0}(\widetilde{\pi}u)(t)+S_{0},$ (3.24)
where we have denoted
$S_{0}:=Q(t_{0})-\star_{0}(\widetilde{\pi}u)(t_{0})\in\Omega^{D}M.$ (3.25)
For future use we also define
$\widetilde{S_{0}}:=\star_{D}S_{0}\in\Omega^{0}M=\star_{D}Q(t_{0})-(\widetilde{\pi}u)(t_{0}).$ (3.26)
@@
Phenomenon Amount $\mathsf{X}]$ Potential $\mathsf{Y}]$ Volumetric capacity $\mathsf{X}\mathsf{Y}^{-1}\mathsf{L}^{-D}]$ Conductivity $\mathsf{X}\mathsf{Y}^{-1}\mathsf{L}^{2-D}\mathsf{T}^{-1}]$ Constitutive law
Mass transfer Mass $\mathsf{M}]$ Mass density $\mathsf{M}\mathsf{L}^{-D}]$ 1 $[1]$ Mass diffusivity $\mathsf{L}^{2}\mathsf{T}^{-1}]$ Fick’s law
Heat transfer Heat energy $\mathsf{M}\mathsf{L}^{2}\mathsf{T}^{-2}]$ Temperature $\theta]$ Volumetric heat capacity $\mathsf{M}\mathsf{L}^{2-D}\mathsf{T}^{-2}\theta^{-1}]$ Thermal conductivity $\mathsf{M}\mathsf{L}^{4-D}\mathsf{T}^{-3}\theta^{-1}]$ Fourier’s law
Charge transport Charge $\mathsf{C}]$ Electric potential $\mathsf{M}\mathsf{L}^{2}\mathsf{T}^{-2}\mathsf{C}^{-1}]$ Volumetric capacitance $\mathsf{M}^{-1}\mathsf{L}^{-2-D}\mathsf{T}^{2}\mathsf{C}^{2}]$ Electrical conductivity $\mathsf{M}^{-1}\mathsf{L}^{-D}\mathsf{T}\mathsf{C}^{2}]$ Ohm’s law
Fluid flow through porous media Volume $\mathsf{L}^{D}]$ Pressure $\mathsf{M}\mathsf{L}^{2-D}\mathsf{T}^{-2}]$ Compressibility $\mathsf{M}^{-1}\mathsf{L}^{D-2}\mathsf{T}^{2}]$ Hydraulic conductivity $\mathsf{M}^{-1}\mathsf{L}^{D}\mathsf{T}]$ Darcy’s law
@@
Discussion 3.18 (Initial conditions).
Since we are describing a transient phenomenon, we need initial conditions. We assume that the initial potential is prescribed to some initial value
@@
Similarly, if advection is present, we will need to know the initial amount
$Q_{0}[\mathsf{Y}]\in\Omega^{D}M,$ (3.41)
with the initial condition
$Q_{0}=Q(t_{0}).$ (3.42)
@@
$u(t_{0})$ $u_{0}$ $\mathsf{Y}$ Initial condition for potential Equation 3.40
$Q(t_{0})$ $Q_{0}$ $\mathsf{X}$ Initial condition for amount Equation 3.42
@@ 3.1 Primal weak formulations of transport phenomena
$\begin{split}\int_{M}w\wedge\frac{\partial Q}{\partial t}&=-\int_{M}(w\wedge d_{D-1}q)+\int_{M}(w\wedge f)\\ &=-\int_{\partial M}\mathop{\rm tr}\nolimits_{\partial M,D-1}(w\wedge q)+\int_{M}(d_{0}w\wedge q)+\int_{M}(w\wedge f)\\ &=-\int_{\Gamma_{\mathbf{N}}}(\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{N}},0}w\wedge\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{N}},D-1}q)+\int_{M}(d_{0}w\wedge(-\star_{1}\widetilde{\kappa}d_{0}u+\star_{D}Q\wedge v))+\int_{M}(w\wedge f)\\ &=-\int_{\Gamma_{\mathbf{N}}}(\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{N}},0}w\wedge g_{\mathbf{N}})-\langle d_{0}w,\widetilde{\kappa}d_{0}u\rangle_{M,1}+\int_{M}(d_{0}w\wedge((\widetilde{S_{0}}+\widetilde{\pi}u)\wedge v))+\int_{M}(w\wedge f)\\ &=-\int_{\Gamma_{\mathbf{N}}}(\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{N}},0}w\wedge g_{\mathbf{N}})-\langle d_{0}w,\widetilde{\kappa}d_{0}u\rangle_{M,1}+\int_{M}(d_{0}w\wedge(\widetilde{S_{0}}\wedge v))+\int_{M}(d_{0}w\wedge(\widetilde{\pi}u\wedge v))+\int_{M}(w\wedge f).\end{split}$ (3.44)
@@
Equating the right-hand sides of Equation 3.44 and Equation 3.45, imposing the Dirichlet boundary condition (Equation 3.19) and the initial conditions (Equation 3.40 and Equation 3.42), leads to the primal weak formulation presented below.
@@
$F_{1}\colon\Omega^{0}M\to\mathbb{R},\quad$ $F_{1}(w):=\int_{M}(w\wedge f)\qquad$ $[\mathsf{X}\mathsf{T}^{-1}],$ (3.46e)
$F_{2}\colon\Omega^{0}M\to\mathbb{R},\quad$ $F_{2}(w):=\int_{M}(d_{0}w\wedge(\widetilde{S_{0}}\wedge v))\qquad$ $[\mathsf{X}\mathsf{T}^{-1}].$ (3.46f)
@@
$\forall w[\mathsf{Y}]\in\mathop{\rm Ker}\nolimits\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{D}},0},\quad$ $B(w,\frac{\partial u}{\partial t})+A_{\mathcal{D}}(w,u)-A_{\mathcal{A}}(w,u)$ $=F_{1}(w)+F_{2}(w)-G(w)\qquad$ $[\mathsf{X}\mathsf{T}^{-1}\mathsf{Y}],$ (3.47a)
@@
$q(t,x)=\begin{cases}(-\star_{1}\widetilde{\kappa}d_{0}u+(\widetilde{S_{0}}+\widetilde{\pi}u)\wedge v)(t,x),&x\notin\Gamma_{\mathbf{N}}\\ g_{\mathbf{N}}(t,x),&x\in\Gamma_{\mathbf{N}}\end{cases},\ t\in I.$ (3.48)
@@
$\forall w[\mathsf{Y}]\in\mathop{\rm Ker}\nolimits\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{D}},0},\quad$ $A_{\mathcal{D}}(w,u)-A_{\mathcal{A}}(w,u)$ $=F_{1}(w)+F_{2}(w)-G(w)\qquad$ $[\mathsf{X}\mathsf{T}^{-1}\mathsf{Y}],$ (3.49a)
@@
$q(x)=\begin{cases}(-\star_{1}\widetilde{\kappa}d_{0}u+\widetilde{Q_{0}}\wedge v)(x),&x\notin\Gamma_{\mathbf{N}}\\ g_{\mathbf{N}}(x),&x\in\Gamma_{\mathbf{N}}\end{cases}.$ (3.50)
@@ 3.2 Mixed weak formulations of transport phenomena
$\begin{split}\langle r,\kappa^{-1}q\rangle_{M,D-1}&=\langle r,\kappa^{-1}q_{\mathcal{D}}\rangle_{M,D-1}+\langle r,\kappa^{-1}q_{\mathcal{A}}\rangle_{M,D-1}\\ &=-\int_{\Gamma_{\mathbf{D}}}(\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{D}},D-1}r\wedge g_{\mathbf{D}})+\langle d_{D-1}r,\tilde{u}\rangle+\langle r,\kappa^{-1}(\star_{D}(S_{0}+\pi\widetilde{u})\wedge v)\rangle_{M,D-1}\\ &=-\int_{\Gamma_{\mathbf{D}}}(\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{D}},D-1}r\wedge g_{\mathbf{D}})+\langle d_{D-1}r,\tilde{u}\rangle+\langle r,\kappa^{-1}(\widetilde{S_{0}}\wedge v)\rangle_{M,D-1}+\langle r,\kappa^{-1}(\star_{D}\pi\widetilde{u}\wedge v)\rangle_{M,D-1}.\end{split}$ (3.54)
@@
Combining the weak version of the constitutive law (Equation 3.54), the weak version of the conservation law (Equation 3.55), the Neumann boundary condition (Equation 3.15), and the initial conditions (Equation 3.40, Equation 3.42), leads to the formulation presented below.
@@
$A\colon\Omega^{D-1}M\times\mathcal{C}^{\infty}(I,\Omega^{D-1}M)\to\mathbb{R},\,$ $A(r,s):=\langle r,\kappa^{-1}s\rangle_{M,D-1}\,$ $[\mathsf{X}^{-1}\mathsf{T}\mathsf{Y}],$ (3.56a)
$B_{\mathcal{D}}\colon\Omega^{D}M\times\mathcal{C}^{\infty}(I,\Omega^{D-1}M)\to\mathbb{R},\,$ $B_{\mathcal{D}}(\widetilde{w},r):=\langle d_{D-1}r,\widetilde{w}\rangle_{M,D}\,$ $[\mathsf{L}^{-D}],$ (3.56b)
$B_{\mathcal{A}}\colon\Omega^{D}M\times\mathcal{C}^{\infty}(I,\Omega^{D-1}M)\to\mathbb{R},\,$ $B_{\mathcal{A}}(\widetilde{w},r):=\langle r,\kappa^{-1}(\star_{D}\pi\widetilde{w}\wedge v)\rangle_{M,D-1}\,$ $[\mathsf{L}^{-D}],$ (3.56c)
$C\colon\Omega^{D}M\times\mathcal{C}^{\infty}(I,\Omega^{D}M)\to\mathbb{R},\,$ $C(\widetilde{w},\widetilde{u}):=\langle\pi\widetilde{u},\widetilde{w}\rangle_{M,D}\,$ $[\mathsf{X}\mathsf{L}^{-2D}\mathsf{Y}^{-1}],$ (3.56d)
$G_{1}\colon\Omega^{D-1}M\to\mathbb{R},\,$ $G_{1}(r):=\int_{\Gamma_{\mathbf{D}}}(\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{D}},D-1}r\wedge g_{\mathbf{D}})\,$ $[\mathsf{Y}],$ (3.56e)
$G_{2}\colon\Omega^{D-1}M\to\mathbb{R},\,$ $G_{2}(r):=\langle r,\kappa^{-1}(\widetilde{S_{0}}\wedge v)\rangle_{M,D-1}\,$ $[\mathsf{Y}],$ (3.56f)
$F\colon\Omega^{D}M\to\mathbb{R},\,$ $F(\widetilde{w}):=\langle f,\widetilde{w}\rangle_{M,D}\,$ $[\mathsf{X}\mathsf{T}^{-1}\mathsf{L}^{-D}],$ (3.56g)
$\mathtt{flow\_rate}\colon\Omega^{0}M\to\Omega^{D-2}M,\,$ $\mathtt{flow\_rate}(w):=\kappa\,d^{\star}_{D}\star_{0}w+(\widetilde{S_{0}}+\tilde{\pi}w)\wedge v\,$ $[\mathsf{X}\mathsf{T}^{-1}\mathsf{Y}^{-1}].$ (3.56h)